OpenStudy (anonymous):

Write an inequality for the given statement: The product of 5 and a number is at least –25

OpenStudy (debbieg):

If I said "z is at least -25", how would you write that?

OpenStudy (anonymous):

z<-25 with the line under it?

OpenStudy (debbieg):

\(z \le -25\) says that z is "less than or equal to" -25, which doesn't make it "at least" -25, does it? It would mean that z is NO BIGGER THAN -25. You want the opposite: z is AT LEAST AS MUCH AS -25....

OpenStudy (debbieg):

Remember, "at least -25" means that it could be =-25, or it could be MORE than -25, but it CAN'T BE LESS than -25.

OpenStudy (anonymous):

ok so it would be z>-24 with the eaqual line under?

OpenStudy (debbieg):

hmmm..... not -24, why would it be -24? \(z \ge -25\) means that z is AT LEAST -25, right? Do you understand?

OpenStudy (anonymous):

yes thank you

OpenStudy (debbieg):

OK, good. Now just replace the "z" with an expression for "the product of 5 and a number".

OpenStudy (anonymous):

\[5n \ge - 25\] ???